OpenStudy (anonymous):

While standing at the edge of the roof of a building, you throw a stone upward with an initial speed of 6.57 m/s. The stone subsequently falls to the ground, which is 17.9 m below the point where the stone leaves your hand. At what speed does the stone impact the ground? How much time is the stone in the air? Ignore air resistance and take g = 9.81 m/s2. (This is NOT a suggestion to carry out such an experiment!)

OpenStudy (anonymous):

@BAdhi

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I need help finding the elapsed time.

OpenStudy (anonymous):

@matt101

OpenStudy (whpalmer4):

While standing at the edge of the roof of a building, you throw a stone upward with an initial speed of 6.57 m/s. The stone subsequently falls to the ground, which is 17.9 m below the point where the stone leaves your hand. At what speed does the stone impact the ground? How much time is the stone in the air? Ignore air resistance and take g = 9.81 m/s2. (This is NOT a suggestion to carry out such an experiment!) Okay, you can write an equation for the height of the stone as a function of time: \[h(t) = -\frac{1}{2}gt^2 +v_0 t + h_0\]where \(g=9.81 \text{ m/s}^2\), \(v_0 = 6.57 \text{ m/s}\) and \(h_0 = 17.9\text{ m}\) as you can see, at \(t=0\) that gives you a height of \(17.9\) meters, with the ground at a height of \(0\) meters. To find the elapsed time, solve the equation for \(h(t) = 0\) and find the (positive) value of \(t\) that makes it true. I suggest solving symbolically and then plugging in numbers, rather than doing algebra with quantities like \(6.57t\). Once you know the time at which the stone hits the ground, you can find the speed at impact by taking the derivative of the height function with respect to time and evaluating it at that time. Have you had calculus? If not, you can also find it by finding the apex of the flight (hint: the vertex of the parabola), and then using the elapsed time from that point to impact to calculate speed, knowing that the acceleration is constant, so \(v = gt\) where \(t\) is the time spent falling from the peak of the flight.

OpenStudy (whpalmer4):

the function I wrote has 3 terms, the first representing the motion due to gravity (hence the negative sign), the second representing the motion due to the initial velocity (positive), and the third representing the starting position.

OpenStudy (anonymous):

I was able to figure out the impact speed but I'm having issues with finding the time.

OpenStudy (anonymous):

I'm not sure what exactly I'm suppose to do to find the time.

OpenStudy (whpalmer4):

how did you find the impact speed?

OpenStudy (anonymous):

(6.57 m/s) +2(-9.81 m/s^2)(17.9)

OpenStudy (anonymous):

*(6.57 m/s)^2

OpenStudy (whpalmer4):

and that equals?

OpenStudy (anonymous):

20.0 because you also take the square root.

OpenStudy (whpalmer4):

uh, square root of a negative quantity?

OpenStudy (anonymous):

That was a typo. *9.81

OpenStudy (anonymous):

Sorry.

OpenStudy (whpalmer4):

okay, I agree that you get the correct answer if you do that, but it isn't 20.0, even if rounded...

OpenStudy (anonymous):

Would it be the 19.86 then?

OpenStudy (whpalmer4):

anyhow, formula for height as a function of time is \[h(t) = -\frac{1}{2}gt^2 + v_0 t + h_0\]as I mentioned, with \(g=9.81,\ v_0=6.57,\ h_0=19.7\) (units omitted) Time of impact is the value of \(t\) which makes \(h(t)=0\). Yes, 19.86 is better than 20.0.

OpenStudy (whpalmer4):

You can use the quadratic formula to solve the equation: \[a x^2 + b x + c = 0\]solutions are \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] You have \(a = -\frac{1}{2}g, \ b = v_0, \ c = h_0\)

OpenStudy (whpalmer4):

sorry, 17.9, not 19.7!

OpenStudy (anonymous):

=-6.57+/-\[\sqrt{6.57^2-4(-4.905)(17.9}\]

OpenStudy (whpalmer4):

all of that divided by -4.905

OpenStudy (anonymous):

right

OpenStudy (anonymous):

13.29

OpenStudy (anonymous):

13.29/-4.905 = -2.71

OpenStudy (whpalmer4):

actually, forgot the \(2\) in \(2a\) for the denominator, so you divide by \(-9.81\)

OpenStudy (whpalmer4):

you want the positive solution, so you need to take the negative value of the numerator

OpenStudy (anonymous):

1.35

OpenStudy (whpalmer4):

no... \[\frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-6.57\pm\sqrt{(6.57)^2-4(-\frac{1}{2})(9.81)(17.9)}}{2*(-\frac{1}{2})(9.81)}\]\[=\frac{-6.57\pm\sqrt{(6.57)^2+2(9.81)(17.9)}}{-9.81}\]\[=\]

OpenStudy (whpalmer4):

Here's a graph of height vs. time:

OpenStudy (anonymous):

Sorry. I keep getting the same answer for some reason.

OpenStudy (whpalmer4):

Well, it's 2.69404 or thereabouts, as you can see on the graph.

OpenStudy (whpalmer4):

Looks like you are dividing by an extra factor of 2 or something.

OpenStudy (anonymous):

When I add the -6.57 I get 1.35 and get -13.29 when I subtract it.

OpenStudy (whpalmer4):

Expand the polynomial before you try to solve it. \[0 = -\frac{1}{2}gt^2 +v_0 t + h_0\]\[0 = -(0.5)(9.81)t^2 + 6.57 t + 17.9\]\[0 = -4.905t^2 + 6.57t + 17.9\]\[a=-4.905, b =6.57, c = 17.9\]\[t = \frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-6.57\pm\sqrt{(-6.57)^2 -4(-4.905)(17.9)}}{2(-4.905)}\]\[=\frac{-6.57-\sqrt{43.1649+351.198}}{-9.81}\]

OpenStudy (anonymous):

I got it this time. I was just entering it into my calculator wrong. Sorry about that.

OpenStudy (anonymous):

Thanks for helping. Now on to the next problems.

OpenStudy (whpalmer4):

Okay, good luck! I have to go now.

OpenStudy (anonymous):

Alright. Thanks again